# Reward Inside the Model: A Lightweight Hidden-State Reward Model for LLM’s Best-of-N sampling

### v1

@@ 4.2 Accuracy
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 42.6 48.2 46.8 50.8 52.0 52.2 54.6 56.8 55.0 51.0 Skywork-Llama3.1-8B 43.8 48.4 48.8 52.2 52.4 53.4 56.8 59.0 61.6 52.9 Starling-7B 39.6 41.2 39.8 50.4 49.0 49.0 53.8 50.2 47.0 46.7 Ultra-13B 44.6 47.4 44.4 53.0 50.6 50.4 53.6 53.0 54.0 50.1 Fine-tuned Eurus-7B 42.8 44.6 45.8 52.0 52.8 50.4 54.2 57.4 57.0 50.8 ELHSR (ours) 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
@@
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 53.0 BoN@1: 80.4 BoN@1: 79.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 69.8 76.2 82.4 87.0 89.0 90.4 87.0 90.8 90.4 84.8 Skywork-Llama3.1-8B 77.2 81.8 85.2 89.6 89.4 89.2 83.0 85.8 87.4 85.4 Starling-7B 62.4 65.6 71.4 83.0 87.2 85.4 78.4 81.2 80.0 77.2 Ultra-13B 72.6 78.6 82.4 88.0 88.6 86.8 82.6 84.6 82.0 82.9 Fine-tuned Eurus-7B 79.0 84.4 86.6 90.0 90.6 90.6 88.2 91.6 92.4 88.2 ELHSR (ours) 80.0 86.4 87.6 90.6 89.6 90.6 89.8 93.2 93.4 89.0
@@
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 42.0 BoN@1: 53.6 BoN@1: 56.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B 44.6 42.8 44.6 58.6 60.6 56.8 48.4 34.6 27.2 46.5 Skywork-Llama3.1-8B 60.4 68.0 69.4 69.0 77.0 79.2 65.6 71.2 72.0 70.2 Starling-7B 51.4 55.0 56.6 64.2 70.8 69.8 56.4 48.6 42.8 57.3 Ultra-13B 58.4 61.6 64.8 67.4 72.8 71.8 58.4 51.4 48.0 61.6 Fine-tuned Eurus-7B 57.6 63.0 57.8 67.8 72.6 70.2 70.8 71.6 70.8 66.9 ELHSR (ours) 61.0 70.4 70.8 69.6 77.0 77.0 75.2 75.8 78.0 72.8
@@
Table 1 reports the Best-of-N performance of each reward model on the MATH, GSM8K and AQuA_RAT datasets. Our ELHSR model consistently outperforms all baseline reward models across most BoN settings, demonstrating the effectiveness of the proposed linear architecture and hinge-loss training objective. A notable observation is that while many baseline reward models exhibit strong performance on specific subsets of the data, ELHSR demonstrates consistent and robust performance across the entire dataset. While the fine-tuned EurusRM-7B delivers respectable results, it still lags behind ELHSR in accuracy and incurs substantially higher computational costs during both training and inference. In contrast, ELHSR not only achieves the best performance but does so with a parameter footprint less than 0.005% that of the baselines and is trained on only 6,000 samples per dataset (see Section 4.3 for detailed comparison). This underscores the exceptional efficiency of our approach.
@@
It is noteworthy that the results presented here were achieved without extensive hyperparameter tuning. Further performance gains could likely be realized through more refined optimization strategies. For example, exploring alternative loss functions, as demonstrated in Appendix Table 6, increasing the training dataset size based on the scalability analysis in Section 4.4, or training ELHSR on a subset of layers as explored in Section 4.5 (Table 3), could yield further improvements. This suggests that ELHSR has significant potential for further optimization and improvement.
Additionally, we further assessed ELHSR’s performance on code comprehension tasks using the CRUXEval-O dataset (Gu et al., 2024a) to broaden our evaluation scope. We use only 600 instances for training for this dataset. The results, detailed in Appendix D (Appendix Table 5), demonstrate that ELHSR outperforms the baselines on this task as well.
4.3 Efficiency Analysis
@@
Figure 3: Comparison of average time and FLOPs per sample for ELHSR and baselines, averaged across different datasets and task-performing models with 1-sigma error bars using standard error of the mean. The y-axis is plotted in log-scale. The results show that ELHSR achieves orders-of-magnitude higher efficiency than the baselines. See Appendix Figure 6 for full details.
@@
ELHSR (on Llama-3.2-3B) $1.8\times 10^{5}$ $6.0\times 10^{3}\times 8$
ELHSR (on Llama-3.1-8B) $2.7\times 10^{5}$ $6.0\times 10^{3}\times 8$
ELHSR (on Ministral-8B) $3.0\times 10^{5}$ $6.0\times 10^{3}\times 8$
Table 2: Comparison of Reward Model Parameters and Training Data Size (number of questions $\times$ number of answers). Our method is remarkably efficient in terms of parameters and data.
We demonstrate ELHSR’s significant parameter efficiency and data efficiency in Table 2. Moreover, we investigate the computational efficiency of ELHSR by comparing the average time and FLOPs required to assign a reward to each sample. Figure 3 presents a comparative analysis of ELHSR and baseline reward models across various datasets and task-performing models. The results reveal that ELHSR achieves orders-of-magnitude improvements in both time and FLOPs compared to the baselines. It is noteworthy that as shown in Section 4.4, even a much smaller training dataset leads to satisfactory performance. This substantial reduction in computational cost, combined with its parameter efficiency and data efficiency, makes ELHSR a highly practical solution for reward modeling in both resource-constrained environments and large-scale applications.
4.4 Scalability with Training Set Size and Reasoning Paths
@@
Figure 4: ELHSR has positive scaling with increased number of training samples and with the number of reasoning paths for inference. The result is averaged across datasets and task-performing models. See Appendix Figure 7 for full results.
@@
Scaling large language models (LLMs) can be achieved through various strategies, broadly categorized as training-time scaling and test-time scaling (Snell et al., 2025; Kaplan et al., 2020; Liu et al., 2025a, b). In this section, we investigate the scalability of ELHSR with respect to both training set size and the number of reasoning paths used during inference.
@@
From Figure 4, our results demonstrate that ELHSR’s accuracy consistently improves with both the number of training samples and the size of the reasoning paths set. Specifically, under different BoN@k settings, we find that ELHSR’s accuracy increases monotonically with training set size. Likewise, for any given training set size, ELHSR’s accuracy also improves as the number of reasoning paths used in the BoN sampling process increases.
@@
It is interesting to note that, as illustrated in Figure 4, the performance of ELHSR with training-time scaling has not appeared to saturate even when there are 6000 training samples, which is the one used in the experiments of Section 4.2. This suggests that further increasing the training set size may lead to even greater performance improvements, highlighting the scalability of ELHSR. This indicates that ELHSR is a compelling way to scale reward modeling and hence is a practical solution for real-world applications.
@@
4.5 Maintaining Performance with Selective Layer Utilization for Higher Efficiency
@@
Table 3: ELHSR’s high Best-of-N sampling performance is maintained even when trained only on layers 16, 24, 28, and 32 for Llama-3.1-8B, leading to a substantial reduction in parameters and improved efficiency.
@@
ELHSR on partial layers 55.8 59.0 60.0 89.4 90.8 92.0 69.6 76.8 77.4 74.5
ELHSR on all layers 55.4 59.4 62.6 90.6 89.6 90.6 69.6 77.0 77.0 74.6
While ELHSR leverages hidden states from all layers of the LLM, we investigated the potential for further parameter reduction by selectively utilizing hidden states from a subset of layers. Specifically, for the Llama-3.1-8B model, which comprises 32 layers (plus an embedding layer), we explored training ELHSR using only the hidden states from layers 16, 24, 28, and 32, following the setting in Azaria and Mitchell (2023). This configuration reduces the parameter count to approximately 4/33 of the full-layer ELHSR model. The results, presented in Table 3, demonstrate that performance remains comparable to the full-layer ELHSR, with certain settings even showing improved accuracy. Notably, on the GSM8K dataset, the BoN@16 and BoN@64 accuracies surpass those of the full-layer ELHSR. As shown in Table 1, the performance of the layer-selected ELHSR also consistently exceeds that of baselines with significantly larger parameter counts, further underscoring the efficiency and effectiveness of our approach.
4.6 Logit-Based Training for Closed-Source LLMs
Table 4: Training ELHSR solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency. For results on GSM8K and AQuA_RAT dataset, please refer to Appendix Table 9.

### v3

@@ 5.2. Accuracy
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 39.0 BoN@1: 47.2 BoN@1: 51.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 42.6 48.2 46.8 50.8 52.0 52.2 54.6 56.8 55.0 51.0 Skywork-Llama3.1-8B (Liu et al., 2024) 43.8 48.4 48.8 52.2 52.4 53.4 56.8 59.0 61.6 52.9 Starling-7B (Zhu et al., 2024) 39.6 41.2 39.8 50.4 49.0 49.0 53.8 50.2 47.0 46.7 Ultra-13B (Cui et al., 2024) 44.6 47.4 44.4 53.0 50.6 50.4 53.6 53.0 54.0 50.1 RLHFlow-8B-Deepseek (Xiong et al., 2024) 43.2 46.4 47.6 51.2 49.6 49.8 56.2 58.4 57.8 51.1 Math-Shepherd-7B (Wang et al., 2024) 43.8 42.6 43.6 48.8 50.8 49.0 55.8 58.8 54.8 49.8 Fine-tuned Eurus-7B 42.2 47.0 46.6 51.8 53.0 50.2 54.8 57.2 57.2 51.1 SWIFT (ours) 49.8 54.6 53.6 55.4 59.4 62.6 57.8 61.6 62.8 57.5
@@
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 53.0 BoN@1: 80.4 BoN@1: 79.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 69.8 76.2 82.4 87.0 89.0 90.4 87.0 90.8 90.4 84.8 Skywork-Llama3.1-8B (Liu et al., 2024) 77.2 81.8 85.2 89.6 89.4 89.2 83.0 85.8 87.4 85.4 Starling-7B (Zhu et al., 2024) 62.4 65.6 71.4 83.0 87.2 85.4 78.4 81.2 80.0 77.2 Ultra-13B (Cui et al., 2024) 72.6 78.6 82.4 88.0 88.6 86.8 82.6 84.6 82.0 82.9 RLHFlow-8B-Deepseek (Xiong et al., 2024) 52.4 53.8 60.0 80.4 77.0 80.0 88.4 91.2 93.4 75.2 Math-Shepherd-7B (Wang et al., 2024) 54.8 49.8 43.4 80.6 75.4 73.0 84.8 86.8 86.8 70.6 Fine-tuned Eurus-7B 79.2 84.4 86.8 90.0 91.2 91.0 88.2 91.8 92.8 88.4 SWIFT (ours) 80.0 86.4 87.6 90.6 89.6 90.6 89.8 93.2 93.4 89.0
@@
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 42.0 BoN@1: 53.6 BoN@1: 56.6 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 44.6 42.8 44.6 58.6 60.6 56.8 48.4 34.6 27.2 46.5 Skywork-Llama3.1-8B (Liu et al., 2024) 60.4 68.0 69.4 69.0 77.0 79.2 65.6 71.2 72.0 70.2 Starling-7B (Zhu et al., 2024) 51.4 55.0 56.6 64.2 70.8 69.8 56.4 48.6 42.8 57.3 Ultra-13B (Cui et al., 2024) 58.4 61.6 64.8 67.4 72.8 71.8 58.4 51.4 48.0 61.6 RLHFlow-8B-Deepseek (Xiong et al., 2024) 53.8 54.0 48.8 63.6 65.0 61.4 64.6 63.8 62.6 59.7 Math-Shepherd-7B (Wang et al., 2024) 52.2 53.4 53.0 64.0 67.8 66.2 63.6 60.4 60.2 60.1 Fine-tuned Eurus-7B 57.4 66.6 62.0 68.8 72.8 69.8 71.2 71.8 71.0 67.9 SWIFT (ours) 61.0 70.4 70.8 69.6 77.0 77.0 75.2 75.8 78.0 72.8
Table 2. Different methods’ best-of-N sampling performance on general domain, including Imbue Code Comprehension, HellaSwag and CoinFlip datasets. Results represent accuracy (percentages omitted). Bold: best performance; Underline: second-best performance; @k: indicates the number of reasoning paths used in best-of-N sampling.
Different methods’ best-of-N sampling performance on general domain, including Imbue Code Comprehension, HellaSwag and CoinFlip datasets.
Imbue Code Comprehension Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 17.2 BoN@1: 47.6 BoN@1: 54.4 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 26.0 28.0 30.2 60.0 62.2 66.2 59.0 61.6 64.0 50.8 Skywork-Llama3.1-8B (Liu et al., 2024) 32.6 47.4 56.0 67.0 70.8 75.2 66.6 71.4 74.8 62.4 Starling-7B (Zhu et al., 2024) 18.8 21.0 21.0 57.6 62.6 64.2 58.6 60.4 62.6 47.4 Ultra-13B (Cui et al., 2024) 28.0 34.8 38.8 61.8 66.6 68.6 60.6 64.4 65.2 54.3 RLHFlow-8B-Deepseek (Xiong et al., 2024) 29.2 37.0 44.4 66.4 71.4 72.0 68.4 74.0 75.2 59.8 Math-Shepherd-7B (Wang et al., 2024) 22.6 25.4 25.4 60.6 60.0 55.8 63.0 65.4 61.4 48.8 Fine-tuned Eurus-7B 35.6 54.4 60.4 67.0 69.4 76.0 67.8 71.0 71.8 63.7 SWIFT (ours) 36.8 57.6 65.6 74.0 84.0 85.0 74.6 82.4 85.4 71.7
HellaSwag Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 55.2 BoN@1: 61.2 BoN@1: 68.2 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 67.4 71.6 72.0 72.6 73.6 74.0 77.8 77.4 78.0 73.8 Skywork-Llama3.1-8B (Liu et al., 2024) 63.0 64.0 67.0 70.6 71.6 71.8 73.0 74.0 76.4 70.2 Starling-7B (Zhu et al., 2024) 53.6 56.2 59.0 64.8 65.8 63.6 66.0 65.0 63.6 62.0 Ultra-13B (Cui et al., 2024) 60.8 64.0 64.2 69.6 69.4 68.2 67.8 67.8 70.6 66.9 RLHFlow-8B-Deepseek (Xiong et al., 2024) 65.2 68.4 69.2 68.2 70.8 71.0 75.2 78.2 79.8 71.8 Math-Shepherd-7B (Wang et al., 2024) 59.8 61.0 62.2 66.0 65.2 63.8 70.2 69.8 73.0 65.7 Fine-tuned Eurus-7B 70.4 76.2 77.4 75.2 79.4 78.6 80.4 81.4 81.6 77.8 SWIFT (ours) 71.0 75.8 77.4 78.0 82.4 83.2 83.4 88.0 88.2 80.8
CoinFlip Dataset
Reward Model Llama-3.2-3B Llama-3.1-8B Ministral-8B Avg. BoN@1: 56.2 BoN@1: 62.6 BoN@1: 20.0 @4 @16 @64 @4 @16 @64 @4 @16 @64 Eurus-7B (Yuan et al., 2025) 65.6 64.2 64.6 83.6 85.6 91.8 45.8 69.8 77.4 72.0 Skywork-Llama3.1-8B (Liu et al., 2024) 72.6 81.0 88.0 88.0 94.4 97.4 25.2 28.2 26.0 66.8 Starling-7B (Zhu et al., 2024) 66.6 73.0 76.2 79.0 84.2 87.2 22.0 19.6 21.0 58.8 Ultra-13B (Cui et al., 2024) 61.6 62.0 59.4 82.0 85.6 84.0 14.0 11.4 11.4 52.4 RLHFlow-8B-Deepseek (Xiong et al., 2024) 64.0 65.6 75.6 83.4 96.4 99.4 34.6 49.4 72.8 71.2 Math-Shepherd-7B (Wang et al., 2024) 55.8 46.4 35.4 62.4 40.4 15.2 20.0 19.0 15.4 34.4 Fine-tuned Eurus-7B 91.2 99.4 99.8 95.2 99.8 100.0 57.8 95.8 99.8 93.2 SWIFT (ours) 91.6 99.8 100.0 95.4 99.6 100.0 57.8 96.2 99.6 93.3
@@
Table 1 reports the Best-of-N performance on mathematics domain of each reward model on the MATH, GSM8K, and AQuA_RAT datasets, and Table 2 reports their performance on other general domains, including code comprehension (Imbue Code Comprehension dataset), commonsense (HellaSwag dataset) and symbolic reasoning (CoinFlip dataset). Our SWIFT model outperforms all baseline reward models across most BoN settings. For example, SWIFT is 12.7% more accurate on average than EurusRM-7B on the MATH dataset, with EurusRM-7B being a state-of-the-art open-source reward model on the math and code domains (Yuan et al., 2025). A notable observation is that while many baseline reward models perform well on specific subsets of the data, SWIFT demonstrates consistent and robust performance across the entire dataset. Although the fine-tuned EurusRM-7B achieves respectable results, it still falls short of SWIFT in accuracy and incurs substantially higher computational costs during both training and inference. In contrast, SWIFT not only achieves the best overall performance but does so with a parameter footprint of less than 0.005% of the baselines and is trained on only 6,000 samples per dataset (see Section 5.3 for detailed comparison). This highlights the exceptional efficiency of our approach.
@@
It is noteworthy that the results presented here were achieved without extensive hyperparameter tuning. Further performance gains could likely be realized through more refined optimization strategies. For example, exploring alternative loss functions, as demonstrated in Appendix Table 7, increasing the training dataset size based on the scalability analysis in Section 5.6, or training SWIFT on a subset of layers as explored in Section 5.7 (Table 5), could yield further improvements. This suggests that SWIFT has significant potential for further optimization and improvement.
5.3. Efficiency Analysis
@@
Figure 3. Comparison of average time and FLOPs per sample for SWIFT and baselines, averaged across different datasets and task-performing models with 1-sigma error bars using standard error of the mean. The y-axis is plotted in log-scale. It shows that SWIFT achieves orders-of-magnitude higher efficiency than baselines. See Appendix Figure 6 and Figure 7 for full details.
Comparison of average time and FLOPs per sample for SWIFT and baselines, averaged across different datasets and task-performing models with 1-sigma error bars using standard error of the mean. The y-axis is plotted in log-scale.
We demonstrate SWIFT’s significant parameter efficiency and data efficiency in Table 3. Moreover, we evaluate the computational efficiency of SWIFT by comparing the average time and FLOPs required to assign a reward to each sample. Figure 3 presents a comparative analysis of SWIFT and baseline reward models across various datasets and task-performing models. The results show that SWIFT achieves orders-of-magnitude improvements in both time and FLOPs compared to the baselines. Notably, as shown in Section 5.6, even a significantly smaller training dataset leads to satisfactory performance. This substantial reduction in computational cost, combined with its parameter efficiency and data efficiency, makes SWIFT a highly practical solution for reward modeling in both resource-constrained environments and large-scale applications.
[…]
Table 3. Comparison of Reward Model Parameters and Training Data Size (number of questions $\times$ number of answers). Our method is remarkably efficient in terms of parameters and data.
Different methods’ best-of-N sampling performance on MATH, GSM8K, and AQuA_RAT datasets.
@@
RLHFlow-8B-Deepseek $8.0\times 10^{9}$ $2.5\times 10^{5}\times 1$
Math-Shepherd-7B $7.2\times 10^{9}$ $4.5\times 10^{5}\times 1$
SWIFT (on Llama-3.2-3B) $1.8\times 10^{5}$ $6.0\times 10^{3}\times 8$
SWIFT (on Llama-3.1-8B) $2.7\times 10^{5}$ $6.0\times 10^{3}\times 8$
SWIFT (on Ministral-8B) $3.0\times 10^{5}$ $6.0\times 10^{3}\times 8$
@@
Figure 4. SWIFT has positive scaling with increased number of training samples and with the number of reasoning paths for inference. The result is averaged across datasets and task-performing models.
SWIFT has positive scaling with increased number of training samples and with the number of reasoning paths for inference.
5.4. Generalization Test
To assess transferability, we pretrain SWIFT on 10k samples of DeepScaleR dataset (Luo et al., 2025) and evaluate it directly on downstream benchmarks (no dataset-specific fine-tuning). With Ministral-8B and BoN@64, SWIFT achieves 62.8/93.6/75.8 on MATH/GSM8K/AQuA_RAT, consistently exceeding all baselines (e.g., fine-tuned Eurus: 57.2/92.8/71.0; Skywork: 61.6/87.4/72.0) and achieving performance comparable to fine-tuned SWIFT. In addition, to further validate that the strong performance of SWIFT does not benefit from fine-tuning solely, we fine-tune Skywork-Llama3.1-8B (the strongest baseline RM on MATH) and evaluate BoN@64 on MATH. Across task LLMs Llama3-3B/Llama3-8B/Ministral-8B, fine-tuned Skywork yields 52.2/54.0/62.6, while SWIFT achieves 53.6/62.6/62.8, consistently outperforming the fine-tuned baseline. This demonstrates that SWIFT’s improvements stem from its design rather than fine-tuning alone, and confirms robust cross-dataset generalization.
5.5. Domains beyond reasoning accuracy (helpfulness/safety)
Table 4. Alignment performance on the PKU-SafeRLHF dataset evaluated by GPT-4o. We report BoN@4 accuracy using Ministral-8B as the policy. SWIFT consistently outperforms strong baselines on both helpfulness and safety dimensions.
Dimension SWIFT Eurus Skywork Ultra RLHFlow Starling Shepherd
Helpfulness 64.2 59.8 63.2 54.6 61.0 60.4 53.0
Safety 92.2 89.8 91.6 79.0 90.2 91.6 84.2
Finally, to test whether SWIFT extends to alignment-oriented evaluation, we run Best-of-$N$ on PKU-SafeRLHF dataset (Ji et al., 2025) using GPT-4o as judge (Ministral-8B; BoN@4). The result is characterized in Table 4. SWIFT significantly outperforms all large text-based reward models. The consistent helpfulness and safety improvement over these baselines suggest that SWIFT’s hidden-state representations capture nuanced response quality beyond reasoning correctness, making it suitable for general alignment tasks.
5.6. Scalability with Training Set Size and Reasoning Paths
@@
Scaling large language models (LLMs) can be achieved through various strategies, broadly categorized as training-time scaling and test-time scaling (Snell et al., 2025; Kaplan et al., 2020). In this section, we investigate the scalability of SWIFT with respect to both training set size and the number of reasoning paths used during inference.
@@
From Figure 4, our results demonstrate that SWIFT’s accuracy consistently improves with both the number of training samples and the size of the reasoning paths set. Specifically, under different BoN@k settings, we find that SWIFT’s accuracy increases monotonically with training set size. Likewise, for any given training set size, SWIFT’s accuracy also improves as the number of reasoning paths used in the BoN sampling process increases.
@@
It is interesting to note that, as shown in Figure 4, the performance of SWIFT with training-time scaling does not appear to plateau even with 6000 training samples, which is the one used in the experiments of Section 5.2. This suggests that further increasing the training set size may yield additional performance improvements, underscoring the scalability of SWIFT. These results highlight SWIFT as a compelling and practical approach to scaling reward modeling for real-world applications.
@@
5.7. Maintaining Performance with Selective Layer Utilization for Higher Efficiency
@@
Table 5. SWIFT’s high Best-of-N sampling performance is maintained even when trained only on layers 16, 24, 28, and 32 for Llama-3.1-8B, leading to a substantial reduction in parameters and improved efficiency.
@@
SWIFT on partial layers 55.8 59.0 60.0 89.4 90.8 92.0 69.6 76.8 77.4 74.5
SWIFT on all layers 55.4 59.4 62.6 90.6 89.6 90.6 69.6 77.0 77.0 74.6
While SWIFT leverages hidden states from all layers of the LLM, we investigated the potential for further parameter reduction by selectively using hidden states from only a subset of layers. Specifically, for the Llama-3.1-8B model, which comprises 32 layers (plus an embedding layer), we explored training SWIFT using only the hidden states from layers 16, 24, 28, and 32, following the setting in Azaria and Mitchell (2023). This configuration reduces the parameter count to approximately 4/33 of the full-layer SWIFT model. The results, presented in Table 5, demonstrate that performance remains comparable to the full-layer SWIFT, with certain settings even showing improved accuracy. Notably, on the GSM8K dataset, the BoN@16 and BoN@64 accuracies surpass those of the full-layer SWIFT. As shown in Table 1, the performance of the layer-selected SWIFT also consistently exceeds that of baselines with significantly larger parameter counts, further underscoring the efficiency and effectiveness of our approach.
We further find that later layers encode stronger correctness signals: using only the first 4 layers gives BoN@64 accuracy 38.2/44.6/45.4 (Llama3-3B/Llama3-8B/Ministral-8B on MATH), while using only the last 4 layers improves it to 55.4/60.4/61.8. Please refer to Appendix E.4 for details.
5.8. Logit-Based Training for Closed-Source LLMs
Table 6. Training SWIFT solely on logits can also yield high performance and can outperform many baselines with orders-of-magnitude higher efficiency. For results on GSM8K and AQuA_RAT dataset, please refer to Appendix Table 12.